OpenStudy (anonymous):

x+2y greater than or equal to 0 graph on coordinate plane

OpenStudy (anonymous):

\[x + 2y \ge 0\]\[2y \ge -x\]\[y \ge -\frac{ 1 }{ 2}x\]